OpenStudy (anonymous):

Consider a function f(x,y) = sqrt(20-x2-7y2). Using linear approximation of f near (2,1), find how large the approximated value of f(1.95, 1.08) is from the exact value a) 0.44% b) 0.28% c) 0.33% d) 0.55% e) None of the above

OpenStudy (anonymous):

linear = f'(x_0)(x-x_0) \[f'(x)=\frac{x+7y}{\sqrt{20-x^2-7y^2}}\] can you do the rest?

OpenStudy (anonymous):

sry, forgot - sign infront of fraction

OpenStudy (anonymous):

No please

OpenStudy (stamp):

\[f(x,y)=\sqrt{20-x^2-7y^2}\]

OpenStudy (stamp):

@myko this is no single variable, this is multivariable

OpenStudy (stamp):

\[f(2,1)=3\]\[f_x(x,y)=-x/\sqrt{20-x^2-7y^2}\]\[f_y(x,y)=-7y/\sqrt{20-x^2-7y^2}\]

OpenStudy (stamp):

\[f_x(2,1)=-2/3\]\[f_y(2,1)=-7/3\]\[L(x,y)=3-\frac{2}{3}(x-2)-\frac{7}{3}(y-1)\]

OpenStudy (stamp):

\[L(1.95,1.08)=2.84667\]\[f(1.95,1.08)=2.83420\]\[2.83420x=2.84667\]\[x=2.84667/2.83420\]\[x=1.00439\] So the linear approximation is ~ 0.44% greater than the actual value